OpenStudy (anonymous):

a jazz concert brought in $135,000 on the sale of 8600 tickets. if the tickets were sold for $10 and $20 each, how many of each type of ticket were sold?

OpenStudy (anonymous):

simultaneous equations 1) \[x + y = 8600\] 2) 10x + 20y = 135000 solve for either x or y and substitute into other equation

OpenStudy (anonymous):

10x + 20y = 135,000 x+y=8600 y= 8600 - x 10x + 20 (8600 - x)= 135, 000 10x +172,000 - 20x = 135,000 -10x + 172,000= 135,000 -10x = -37000 x= 3700 ($10 tickets) 3700 + y = 86000 y= 82,300 (20 tickets)

OpenStudy (anonymous):

let's see x = 8600 - y so 10(8600-y) + 20y = 135000 86000 -10y +20y = 135000 10y = 49000 y = 4900 x = 8600 - 4900 = 3700

OpenStudy (anonymous):

@ nicole you got one to many zeros in your last equation